what is the common solution for the three linear equation x-y=4,x+y=6,x-5y=0
Answer by Alan3354(69443)

what is the common solution for the three linear equation x-y=4,x+y=6,x-5y=0
-------------
There might not be one. If you graph 3 lines, they don't necessarily have an intersection of all 3.
---------------
x-y=4
x+y=6
------- Add
2x = 10
x = 5, y = 1
-------------
x - 5y = 0 also fits, so that's the solution.